There are 2 firms in a market producing differentiated products. The firms both have MC that is equal to 0

Firm 1 demand is
q1(p1,p2) = 6-2p1 + p2

Firm 2 demand is
q2(p1,p2) = 6-2p2 + p1

1. Firms compete in quantities- Cournot Competition. What are the inverse demand functions for firm 1 and 2?

2. Find and graph each firm’s best response functions. The quantities are strategic substitutes or complements?

3. Find the Nash equilibrium prices and quantities when both firms use Cournot. What are the firm’s profits?
Answer #1

Page numbers are given at the top of the page, follow them.

Vi= 6-2P,+ PR. 1 9:56=2P2 +P, >> 28,-P,= 6-9; 230 2P, -P,36-93 2) 3P, -2,36---01:. 4px - 21,5. 12. = 24-- Sebring the above tand TC = 0; an The profit function of at firm is given by; T Total Revenue - Total Cast | A: FR - TC , T = PxQ - o TR = PxQ7from the demand function ginene in the quaetion we can identify the nature of the goods. when P increases there is an increasSetting 9, - 3.6 bine equation © ; . 9,= 4.5 - .0, >>V=4.5 - 5* 3.6 >> 9: 3.6 The equlibrium quantity for the two firms are;UL Profit of firm 1: T;= P, xq, = 3.6* 2:4 = 8.64. Profit of firem 2 - T = P₂ X %= 8.68 2:4=8.64.